a)Evaluate f(5)where f(x)=x2+(12-5x)/8x
F(5) =
b)Evaluate g(x+3) where g(x)=3x2-6x+8.
c)Evaluate h(3x)where h(y)=y3-6/7y.

Answer by Alan3354(69443)

a)Evaluate f(5)where f(x)=x2+(12-5x)/8x
F(5) =
---
Wherever you see an x, sub 5.
----------------------
b)Evaluate g(x+3) where g(x)=3x2-6x+8.
Same as a above.
-----------
c)Evaluate h(3x)where h(y)=y3-6/7y
Same as a above.
=====================
It's a simple matter of substitution, tho I'm not sure what y3 means.